• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 16:11:29 on June10,2000

    Entry number 43373

    keyword=shift summary

    8:15 Run 1890 1.3M events, 68.9 microamps,
    8:25 HRSE Q3 tripped again. End run 1891, Request Hall A access to reset. Mark Stevens here. Q3 was not ramped up to 1600.
    8:55 Mark Stevens reset magnet and cleaned connection in power supply. Waiting for beam.
    9:04 Beam back. Started run 1892
    9:20 Run 1892 ended, 1.3M events, 67.3 ua,
    9:35 Run 1893 ended, 1.3M events, 67.5 ua,
    9:45 Beam off due to MCC FOP checks
    10:00 Beam back
    10:05 HRSE Q3 tripped again. Called MCC. Hall A access., Stevens coming again.
    11:20 Stevens cleaned more switches. Q3 back up via proper procedure. Air condition repair guys still outside on the mound.
    12:25 Beam back
    13:05 MCC retuning beam.
    13:25 Beam back
    13:- Beam off
    14:35 Beam on
    15:15 Beam off due to rash of trips at MCC
    15:45 Beam on





    Start Time.....Run Number....Events.....Current(ua)......Charge(mc)
    8.02................1890.................1.3M.........68.9...................51.8
    8:16................1891.................1.17M.......68.5...................44.4
    9.04................1892.................1.3M.........67.3...................53.6
    9.19................1893.................1.3M.........67.5...................53.8
    9.37................1894.................0.59M.......67.3...................23.6
    10:00..............1895.................0.18...........66.8.................. 8.1
    12:24..............1896.................1.3M.........65.5...................53.5
    12.44..............1897.................1.3M.........65.1...................52.8
    13.04..............1898.................1.3M.........63.1...................54.5
    14:45..............1899.................1.3M.........63.3...................53.5
    15:00..............1900.................1.3M.........61.3...................54.0